find f^-1(x) f(x)= -8x + 48
Found 2 solutions by stanbon, Theo:
Answer by stanbon(57361)

Find f^-1(x) if f(x)= -8x + 48
-----
y = -8x + 48
Interchange x and y to get:
x = -8y + 48
Solve for "y":
-8y = x-48
y = -(x-48)/8
That is f^-1(x)
==========================
Cheers,
Stan H.

Answer by Theo(3464)

I think you want the inverse function of f(x) = -8x + 48

solve for x.

add 8x to both sides of this equation and subtract f(x) from both sides of this equation to get:

8x = -f(x) + 48

divide both sides of this equation by 8 to get:

x = -f(x)/8 + 6

replace x with f^-1(x) and f(x) with x to get:

f^-1(x) = -x/8 + 6

that's your inverse equation.